Indicate the function of f(x) belowin fourier series

0,-0,5<x<0

F(x) ={

x, 0<x<0, 5


1
Expert's answer
2022-01-10T14:06:57-0500
"a_0=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)dx"

"=\dfrac{1}{0.5}\displaystyle\int_{0}^{0.5}xdx=[x^2]\begin{matrix}
 0.5 \\
 0
\end{matrix}=0.25"

"a_n=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)\cos\dfrac{n\pi x}{L} dx"

"=\dfrac{1}{0.5}\displaystyle\int_{0}^{0.5}x\cos\dfrac{n\pi x}{0.5}dx=2\displaystyle\int_{0}^{0.5}x\cos(2\pi nx)dx"

"=2[\dfrac{x\sin(2\pi n x)}{2\pi n}+\dfrac{\cos(2\pi n x)}{4\pi^2n^2}]\begin{matrix}
 0.5\\
0
\end{matrix}"

"=\dfrac{(-1)^n-1}{2\pi^2n^2}=-\dfrac{1}{\pi^2(2k+1)^2}, k\in Z^+"



"b_n=\dfrac{1}{L}\displaystyle\int_{-L}^{L}f(x)\sin\dfrac{n\pi x}{L} dx"

"=\dfrac{1}{0.5}\displaystyle\int_{0}^{0.5}x\sin\dfrac{n\pi x}{0.5}dx=2\displaystyle\int_{0}^{0.5}x\sin(2\pi nx)dx"

"=2[-\dfrac{x\cos(2\pi n x)}{2\pi n}-\dfrac{\sin(2\pi n x)}{4\pi^2n^2}]\begin{matrix}
 0.5\\
0
\end{matrix}"

"=\dfrac{(-1)^n}{2\pi n}=-\dfrac{(-1)^k}{2\pi (k+1)}, k\in Z^+"

"f(x)=\dfrac{1}{8}-\displaystyle\sum_{k=0}^{\infin}(\dfrac{\cos(2\pi (2k+1) x)}{\pi^2(2k+1)^2}+\dfrac{(-1)^k\sin(2\pi (k+1) x)}{2\pi (k+1)})"